Ten years ago, Lucas Inc. earning $0.45 per share.  Its earnings this year were $2.28. What was the growth rate in earnings per share (EPS) over the 10-year period?

15.17% / 15.97% / 16.77% / 17.62% / 18.49%
Answer #1

Option (d) is correct

Here we will use the following formula:

FV = PV * (1 + r%)n

where, FV = Future value = $2.28, PV = Present value = $0.45, r = rate of interest, n= time period = 10

now, putting theses values in the above equation, we get,

$2.28 = $0.45 * (1 + r)10

$2.28 / $0.45 = (1 + r)10

5.06667 = (1 + r)10

(5.06667)1/10= 1 + r

1.1761757 = 1 + r

r = 1.1761757 - 1

r = 0.1762 or 17.62%

So, growth rate was 17.62%